A firm has a stock price of $60.00 per share. The firm's earnings are $80 million, and the firm has 25 million shares outstanding. The firm has an ROE of 16% and a plowback of 70%. What is the firm's PEG ratio? 1.54 1.30 1.67 1.17
Answer #1
Growth rate=ROE*retention ratio
growth rate=16*0.7
growth rate = 11.2

PEG = (Price/(earnings/shares))/growth rate

=(60/(80/25))/11.2= 1.67